# Parametric Instabilities of Correlated Quantum Matter

### v1

@@ Abstract
Strongly correlated quantum materials exhibit a rich landscape of ordered phases with highly tunable properties, making them an intriguing platform for exploring non-equilibrium phenomena. A key to many of these phases is collective bosonic excitations, encoding fluctuations of the underlying order. In this work, we develop a general theoretical framework for parametric driving of such modes, whereby periodic modulation of microscopic parameters generates resonant two-boson processes. We show that the feasibility and strength of this drive depend sensitively on whether the targeted parameter alters the properties of the bosonic excitations vacuum, linking potential parametric instabilities directly to the fidelity susceptibility of the ground state. The driving facilitates nonthermal melting of the parent orders, as well as stabilization of novel steady states with experimentally distinct signatures. Through microscopic case-studies of correlated electronic systems, we identify promising driving knobs, highlight the role of quantum geometry in the collective modes susceptibility, and propose realistic experimental probes. Collective excitations are a powerful resource for steering correlated phases out of equilibrium, and will likely have several applications in quantum science. Our work provides the toolbox for controlling these excitations.
@@ I Introduction
Recent years have been marked by the rise of correlated quantum materials, that host a variety of exotic quantum phases and offer extensive in situ tunability. A non-exhaustive list includes moiré graphene or transition metal dichalcogenides (TMDs) materials [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], monolayer TMDs [19, 20, 21], crystalline graphene superconductors and fractional Chern insulators [22, 23, 24, 25, 26, 27, 28, 29, 30], fractional quantum Hall (FQH) states in bilayer graphene [31, 32, 33, 34], excitonic double layers  [35, 36, 37, 38], and tunable oxide interfaces [39, 40, 41]. The combination of a rich strongly-interacting phase diagram with extraordinary tunability unlocks a myriad of unexplored opportunities. Proposals involving the dynamical manipulations of different parameters of the system, whether by a rapid quench [42, 43, 44, 45, 46, 47] or by periodic driving [48, 49, 50] hold promise to deepen our understanding of these materials, and to access and manipulate novel phases of matter.
Most of the ordered correlated phases mentioned above possess collective low-energy excitations, which may be thought of as infinitesimal perturbations to the global order. These modes garnered significant interest, as their properties may help account for some unusual experimental observations, e.g., an entropic Pomeranchuk effect [51], reduction of ferromagnetic Curie temperatures [52], and unconventional superconductivity [53, 54]. So far, however, in the context of the novel systems discussed above, the vast potential of these collective excitations as a resource for out-of-equilibrium physics remains underexplored.
@@
In this work we study parametric driving of the low-lying bosonic collective modes in correlated electronic systems. Such driving may be achieved when the underlying parameters which determine and control the ground-state phase diagram are periodically modulated at a frequency $\omega_{d}$, matching the sum of the frequencies of two bosonic excitations, see Fig. 1a. Such periodic tuning of parameters may be accomplished by various tuning knobs, including electronic density, applied electric displacement field, and even the effective strength of electronic interactions [55, 56, 57]. This modulation simultaneously modifies the soft-mode spectrum. This sort of modification is demonstrated in Fig. 1b, concerning a specific microscopic model which will be later introduced.
Not every modulation, however, leads to an effective parametric drive, as we discuss in Sec. II. Concretely, effective modulations must induce a time-dependent two-boson term in the driven collective Hamiltonian. A meticulous analysis of the microscopic details and dynamical properties of these excitations is necessary in order to determine which modulation will best parametrically drive them. We present a straightforward treatment of the collective excitations atop the correlated phase, directly addressing this subtle yet important issue in Sec. III.
@@
Our work shows that parametric driving of collective modes can only be made by microscopic parameters that alter the vacuum state of the targeted excitations. Consequently, a system’s response to parametric driving at a certain point on the phase diagram probes the nature of quantum fluctuations in the relevant ground-state. Fig. 1c illustrates the relationship between the modulation of a parameter in the microscopic Hamiltonian and the resulting change in the effective Hamiltonian of the low-energy collective fluctuations. It is the induced two-boson term (bottom right panel of Fig. 1c) that corresponds to the strength of the parametric driving. The same term is also related to the fidelity susceptibility of the vacuum of fluctuations, revealing the intimate connection between these two quantities.
The consequences of triggering a parametric instability may be far-reaching. Once the driving strength exceeds the inverse-lifetime of the driven mode, one expects the parent order to melt, albeit in a nonthermal way (cf. Ref. [58]). If the parametric drive is counterbalanced by nonlinear dissipation terms (due to, e.g., interactions between the bosonic modes), however, the driving can result in a non-equilibrium steady-state [59, 60, 61]. Curiously, thoughtful targeting of specific collective excitations may drive the system towards a low-lying order just outside the reach of the equilibrium phase diagram.
In Sec. IV we study detailed microscopic examples of correlated electronic systems, their relevant low-energy excitations, as well as their susceptibility to parametric driving by common experimental controls. These vignettes also shed light on the ubiquitous interplay between quantum geometry and collective modes in the presence of electronic interactions (cf. Ref. [62]). Specifically, it appears that systems are most susceptible to changes of quantum geometric parameters when the interaction range is comparable to the length scale set by quantum geometry. We further relate these microscopic examples to realistic systems, propose which parameters are most promising for driving, the modulation amplitudes necessary, and discuss measurable experimental consequences of the respective instabilities.
We conclude this work in Sec. V, summarizing our findings and shortly discussing further implications of this study. Namely, the role collective excitations in correlated electronic materials play in scenarios involving a quantum quench, and the potential such systems hold for quantum information applications when the low-lying bosonic modes are coupled non-trivially to other systems.
@@ II.1 Harmonic oscillator
$H_{\rm harm.}=\frac{1}{2}\left(m^{-1}\mathbb{p}^{2}+k\mathbb{x}^{2}\right)=\omega\left(a^{\dagger}a+\frac{1}{2}\right),$ (1)
where $\mathbb{x}$ and $\mathbb{p}$ are canonically-conjugate operators, $\left[\mathbb{x},\mathbb{p}\right]=i$, $\omega=\sqrt{m^{-1}k}$, and $a=\frac{1}{\sqrt{2}}\left(\sqrt{m\omega}\mathbb{x}+\frac{i}{\sqrt{m\omega}}\mathbb{p}\right)$ is a bosonic annihilation operator.
How does a small change of the bare parameters in Eq. (1) affect the system? Concretely, modifying $k\to k+\delta k$, $m^{-1}\to m^{-1}+\delta m^{-1}$, brings about a two-boson term $\delta z\left(a^{2}+a^{\dagger 2}\right)$  [^1] Introducing a $\propto\left\{\mathbb{x},\mathbb{p}\right\}$ perturbation will change the relative phase between $a^{2}$ and $a^{\dagger 2}$., with
$\delta z=\frac{1}{2}\omega\left(\frac{\delta k}{k}-\frac{\delta m^{-1}}{m^{-1}}\right).$ (2)
@@
An immediate consequence of Eq. (2) is that parametric driving necessitates changing the ratio $k/\left(m^{-1}\right)$. This ratio is nothing but the oscillator ground-state fluctuation squeezing, i.e.,
$\frac{\langle\mathbb{p}^{2}\rangle_{0}}{\langle\mathbb{x}^{2}\rangle_{0}}=\frac{k}{m^{-1}},$ (5)
@@ II.2 Application of parametric driving to collective modes: Preliminaries
Figure 2: Squeezing of vacuum fluctuations for generic ordered states. (a) Schematic depiction of the Bloch sphere for an SU(2) symmetric or easy-axis (along $\hat{z}$) ferromagnet. The ordering direction is indicated by the black arrow. Small fluctuations about the ordered state are represented by the red cloud. A change in the parameters of the parent Hamiltonian of this phase cannot change the nature of the fluctuations without explicitly breaking the remaining U(1) symmetry. (b) Bloch sphere of an easy-plane ferromagnet. The out-of-plane direction is inherently harder for the vacuum fluctuations as compared to in-plane. The fluctuation vacuum is thus squeezed (red oval cloud). Modification of Hamiltonian parameters will generically alter the amount of the aforementioned squeezing, depicted by the elongated blue cloud. The vacuum of fluctuations is susceptible to such parameter changes. (c) Generalized Bloch sphere for an antiferromagnet. The north pole represents the ordered state. The two fluctuation quadratures correspond to either a rotation of the Néel vector, (an easy direction) or to ferromagnetic alignment (hard direction). Squeezing, sensitive to microscopic Hamiltonian parameters, is thus an unavoidable consequence.
@@ III.1 Derivation of the collective mode Hamiltonian
$\equiv\frac{1}{2}\left(1+{\cal O}_{\mu\nu}\left({\bf k}\right)\right).$ (8)
With no loss of generality, we assume that the density matrix is diagonal in momentum $\bf k$  [^2] The limitations set by this assumption are rather moderate, since most density wave orders may be captured by working in a basis where one assigns an additional “flavor” index to sections of the Brillouin zone before potentially coupling them..
The operator $\cal O$ can be thought of as the order parameter of $|\Psi_{0}\rangle$. The Slater determinant density matrix is a projection at any $\bf k$ point, $P^{2}\left(\bf k\right)=P\left(\bf k\right)$ (its eigenvalues are all 0 or $1$). Thus, ${\cal O}^{2}\left({\bf k}\right)=1$, and one finds $\frac{1}{\Omega}\sum_{\bf k}\left\langle{\cal O}_{\bf k}\right\rangle=n$, where $n\equiv\frac{1}{\Omega}\sum_{\bf k}{\rm Tr}P_{\bf k}$ is the density. Henceforth, we will assume for simplicity the absence of a Fermi surface in the ordered state. As a consequence, both $P$ and $\cal O$ are now $\bf k$-independent matrices. This restriction to insulating or semi-metallic ordered states is not necessary in principle. It may be relaxed, within the same framework presented below, by a more complicated treatment of the fluctuations as compared to the one that follows (cf. Ref. [74]).
@@
The matrix $\phi_{\bf Q}$ consists of the various complex amplitudes that parametrize the perturbation from the approximate ground state. Crucially, in order for $^{\hat{F}}$ to be able to generate fluctuations about the ordered state, the matrix $\phi_{\bf Q}$ must anticommute with the order parameter, $\left\{\phi_{\bf Q},{\cal O}\right\}=0$. Pieces of $\phi_{\bf Q}$ which commute with $\cal O$ would result in the same density matrix for $|\Psi_{\phi}\rangle$ as in Eq. (8), and thus do not generate fluctuations.
@@
$\phi^{\mu\nu}_{{\bf Q}}\left({\bf p}\right)=\sum_{i}\left[{{\cal M}^{i}}\right]^{\mu\nu}\Phi_{{\bf Q}}^{i}\left({\bf p}\right)$ (11)
@@ III.2 Parametric term
[…]
@@
So we may finally extract $\delta{\cal H}$ by the simple expression,

### v2

@@ Abstract
Collective bosonic excitations are a fascinating aspect of broken-symmetry correlated phases. A wealth of such phases emerged in tailored moiré heterostructures, where, in addition, new direct knobs of control exist. Our work explores how the associated collective bosonic modes can be directly manipulated and amplified via parametric driving. As we show, parametric driving is closely related to many-body quantum geometry, as it reveals information about the fidelity susceptibility of the underlying ground state with respect to changes in the control knobs of the interacting system. The parametric response hinges on the tunability of the hierarchy of competing ground states, which we show manifests as squeezing of the bosonic mode vacuum. Furthermore, parametrically-driven modes can produce a high-amplitude modulation in the system that could be easily observed, and could also be used to produce a new non-equilibrium prethermal states with different symmetries. We derive a general framework for parametric amplification of collective bosonic modes in broken-symmetry phases, and provide case-studies of current relevance. Elucidating the connection between the dynamics of these excitations and microscopic electronic observables is key to harnessing their immense potential.
@@ I Introduction
Recent years have been marked by the discovery of correlated quantum materials, that host a variety of exotic quantum phases and offer extensive in situ tunability. A non-exhaustive list includes moiré graphene or transition metal dichalcogenides (TMDs) materials [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], monolayer TMDs [19, 20, 21], crystalline graphene superconductors and fractional Chern insulators [22, 23, 24, 25, 26, 27, 28, 29, 30], fractional quantum Hall (FQH) states in bilayer graphene [31, 32, 33, 34], excitonic double layers  [35, 36, 37, 38], and tunable oxide interfaces [39, 40, 41]. The combination of a rich strongly-interacting phase diagram with extraordinary tunability unlocks a myriad of unexplored opportunities. Proposals involving the dynamical manipulations of different parameters of the system, whether by a rapid quench [42, 43, 44, 45, 46, 47] or by periodic driving [48, 49, 50] hold promise to deepen our understanding of these materials, and to access and manipulate novel phases of matter.
Most ordered correlated phases possess collective low-energy excitations. These modes are often at the center of unusual experimental observations, e.g., an entropic Pomeranchuk effect [51], reduction of ferromagnetic Curie temperatures [52], and unconventional superconductivity [53, 54]. We would like to argue that these collective excitations are also an untapped resource for out-of-equilibrium phenomena in the correlated materials that host them.
Figure 1: Driving parametric instabilities of correlated electrons. We consider periodic modulation of the microscopic parameters of a strongly interacting material, represented schematically by $p\left(t\right)$ (e.g. ac gate-voltage modulation). As a result, the effective low-energy degrees of freedom may experience a collective resonant instability, where the the fluctuations around the correlated ordered state grow exponentially. The latter can eventually saturate, and manifest a non-equilibrium steady state with textured excitations. The response of the system is intimately related to the fidelity susceptibility $\chi_{F}$ of the correlated ground state, which tends to diverge in the vicinity of phase transitions.
@@
In this work we study parametric driving of the low-lying bosonic collective modes in correlated electronic systems. Particularly, we explore the effect of periodic modulation of the control parameters of the system at a frequency $\omega_{d}$, which matches the sum of the frequencies of two bosonic excitations, see Fig. 2a. The control parameters we consider are typical experimental knobs such as electronic density, applied electric displacement field, or the screening of electronic interactions [55, 56, 57]. Such modulations also affect the soft-mode spectrum as demonstrated in Fig. 2b, in the microscopic model which we introduce in IV.
Parametric drives are particularly intriguing since they connect changes of underlying system parameters with an exponentially fast build-up of amplitude. We first discuss how an effective parametric drive is achieved (Sec. II). Effective modulations must induce a time-dependent two-boson Hamiltonian term. Whether such a term emerges, and how strong it is, depend crucially on the microscopic details and dynamical properties of the targeted excitations as well as the suggested knob. In Sec. III we develop a general yet straightforward theoretical framework for deriving the properties of collective modes and their parametric-driving terms. We show how the parametric driving emerges, and connect its strength with the quantum-geometric response of the underlying many-body ground state to the driving knobs.
@@
Our approach reveals that parametric driving can only be mediated by modulations that alter the vacuum state of the targeted excitations. In fact, parametric driving strength is a direct probe of the quantum ground-state fluctuations. Fig. 2c highlights how modulation of a microscopic Hamiltonian parameter translates to the effective Hamiltonian of the low-energy collective fluctuations. It is the induced two-boson term (bottom right panel of Fig. 2c) that corresponds to the strength of the parametric driving. Remarkably, the same term is intimately related to the susceptibility of the vacuum of fluctuations. Thus, efficiently driving and probing the parametric instability of a correlated phase with respect to microscopic parameters is also a probe of the fidelity susceptibility of the ground state. This, in turn, provides an independent mapping of phase boundaries, and may reveal hidden transitions.
Triggering parametric instabilities could have dramatic signatures. Once the parametric response of the system exceeds the inverse-lifetime of the driven mode, one expects the parent order to melt, albeit in a nonthermal way (cf. Ref. [58]). If the parametric drive is counterbalanced by nonlinear dissipation terms (due to, e.g., interactions between the bosonic modes or more intricate sources of non-linearity), however, the driving can result in a non-equilibrium steady-state [59, 60, 61, 62]. Curiously, thoughtful targeting of specific collective excitations may drive the system towards a low-lying order just outside the reach of the equilibrium phase diagram.
Sec. IV concentrate on two examples of current interest: quantum Hall bilayers and moiré graphene. We work out their relevant low-energy excitations, their susceptibility to parametric driving and relate those to the energetical hierarchy of competing many-body ground states. These examples demonstrate the ubiquitous interplay between quantum geometry and collective modes in the presence of substantial electronic interactions (cf. Ref. [63]). Specifically, it appears that systems are most susceptible to changes of quantum geometric parameters when the interaction range is comparable to the length scale set by quantum geometry. We further relate these microscopic examples to realistic systems, propose which parameters are most promising for driving, the modulation amplitudes necessary, and discuss measurable experimental consequences of the respective instabilities.
Sec. V, provides a summary of our results, and discusses further implications of this study. Namely, the role collective excitations in correlated electronic materials play in scenarios involving a quantum quench, and the vast potential such systems hold for quantum information applications when the low-lying bosonic modes are coupled non-trivially to other systems.
@@ II.1 Harmonic oscillator
$H_{\rm harm.}=\frac{1}{2}\left(\frac{1}{m}\mathbb{p}^{2}+k\mathbb{x}^{2}\right)=\omega\left(a^{\dagger}a+\frac{1}{2}\right),$ (1)
where $\mathbb{x}$ and $\mathbb{p}$ are canonically-conjugate operators, $\left[\mathbb{x},\mathbb{p}\right]=i$, $\omega=\sqrt{\frac{k}{m}}$, and $a=\frac{1}{\sqrt{2}}\left(\sqrt{m\omega}\mathbb{x}+\frac{i}{\sqrt{m\omega}}\mathbb{p}\right)$ is a bosonic annihilation operator.
How does a small change of the bare parameters in Eq. (1) affect the system? Concretely, modifying $k\to k+\delta k$, $\frac{1}{m}\to\frac{1}{m}-\frac{\delta m}{m^{2}}$, brings about a two-boson term $\delta z\left(a^{2}+a^{\dagger 2}\right)$  [^1] Introducing a $\propto\left\{\mathbb{x},\mathbb{p}\right\}$ perturbation will change the relative phase between $a^{2}$ and $a^{\dagger 2}$., with
$\delta z=\frac{1}{2}\omega\left(\frac{\delta k}{k}+\frac{\delta m}{m}\right).$ (2)
@@
An immediate consequence of Eq. (2) is that parametric driving necessitates changing the ratio $k/\left(\frac{1}{m}\right)$. This ratio is nothing but the oscillator ground-state fluctuation squeezing (Fig. 3a), i.e.,
$\frac{\langle\mathbb{p}^{2}\rangle_{0}}{\langle\mathbb{x}^{2}\rangle_{0}}=km,$ (5)
@@ II.2 Application of parametric driving to collective modes: Preliminaries
Figure 3: Squeezing of vacuum fluctuations for generic ordered states. (a) For the quantum harmonic oscillator, the ratio of uncertainty of the two quadratures in the ground state depends on parameters of the oscillator, see Eqs. (1),(5). The squeezing in phase space is shown schematically as it evolves with the squeezing parameter $km$. (b) Schematic depiction of the Bloch sphere for an SU(2) symmetric or easy-axis (along $\hat{z}$) ferromagnet. The ordering direction is indicated by the black arrow. Small fluctuations about the ordered state are represented by the red cloud. A change in the parameters of the parent Hamiltonian of this phase cannot change the nature of the fluctuations without explicitly breaking the remaining U(1) symmetry. (c) Bloch sphere of an easy-plane ferromagnet. The out-of-plane direction is inherently harder for the vacuum fluctuations as compared to in-plane. The fluctuation vacuum is thus squeezed (red oval cloud). Modification of Hamiltonian parameters will generically alter the amount of the aforementioned squeezing, depicted by the elongated blue cloud. The vacuum of fluctuations is susceptible to such parameter changes. (d) Generalized Bloch sphere for an antiferromagnet. The north pole represents the ordered state. The two fluctuation quadratures correspond to either a rotation of the Néel vector, (an easy direction) or to ferromagnetic alignment (hard direction). Squeezing, sensitive to microscopic Hamiltonian parameters, is thus an unavoidable consequence.
@@
II.3 Parametric efficiency
Practically speaking, one is interested in the precise strength of the perturbation-induced parametric term. Given the expected exponential growth induced by it, it is reasonable to compare the energy of the two-boson term of interest, $\delta z$, to the energy of the driven collective mode ${\cal E}_{\bar{n},{\bf Q}}$. This is an effective measure of the number of evolution cycles of that mode before the parametric amplification is pronounced, i.e., the exponential prefactor exceeds $O\left(1\right)$.
We define the parametric efficiency $\eta_{{\rm par}.}^{\left(p\right)}$ with respect to a small change $\delta p$ in the control parameter $p$ [in direct analogy to Eq. (2), where the efficiency is unity with respect to $k$ and $m^{-1}$],
In line with the preceding discussion, we find [see Eq. (41)] that the parametric efficiency may exceed unity and diverge as the squeezing of the vacuum fluctuations becomes more pronounced.
We note that linear dissipation would compete with the driving, leading to reduced effective efficiency $\tilde{\eta}_{\rm par.}^{\left(p\right)}=\eta_{\rm par.}^{\left(p\right)}\left(1-\frac{{\cal E}_{\bar{n},{\bf Q}}}{Q\delta z}\right)$, where $Q$ is the quality factor of the driven mode.
@@ III.1 Derivation of the collective mode Hamiltonian
$\equiv\frac{1}{2}\left(\delta_{\mu\nu}+{\cal O}_{\mu\nu}\left({\bf k}\right)\right).$ (9)
With no loss of generality, we assume that the density matrix is diagonal in momentum $\bf k$. Spontaneous formation of density wave orders, which defy this assumption, may be captured by working in a basis where one assigns an additional flavor index to sections of the Brillouin zone before coupling them, i.e., a folded-zone scheme.
The operator $\cal O$ can be thought of as the order parameter of $|\Psi_{0}\rangle$. The Slater determinant density matrix is a projection at any $\bf k$ point, $\left[P\left(\bf k\right)\right]^{2}=P\left(\bf k\right)$ (its eigenvalues are all 0 or $1$). Thus, $\left[{\cal O}\left({\bf k}\right)\right]^{2}=1$, and one finds $\frac{1}{\Omega}\sum_{\bf k}\left\langle{\cal O}\right({\bf k}\left)\right\rangle=n$, where $n\equiv\frac{1}{\Omega}\sum_{\bf k}{\rm Tr}P_{\bf k}$ is the density.
Henceforth, we will assume for simplicity that $P$ (and therefore $\cal O$) are $\bf k$-independent matrices. In some cases, this may entail applying a unitary transformation, e.g., a Bogoliubov transformation, in order to bring $P$ to this desired form. This assumption restricts our present discussion to cases where the ordered state lacks a Fermi surface. This restriction to insulating or semi-metallic ordered states is not necessary in principle. It may be relaxed, within the same framework presented below, by a more complicated treatment of the fluctuations as compared to the one that follows (cf. Ref. [74]).
@@
The matrix $\phi_{\bf Q}$ consists of the various complex amplitudes that parametrize the perturbation away from the approximate ground state. Crucially, pieces of the matrix $\phi_{\bf Q}$ which commute with $\cal O$ would not create a particle-hole excitation of the $|\Psi_{0}\rangle$ state, and lead to $|\Psi_{\phi}\rangle$ possessing the same density matrix as in Eq. (9). Thus, $\phi_{\bf Q}$ anticommutes with the order parameter, $\left\{\phi_{\bf Q},{\cal O}\right\}=0$.
@@
$\phi^{\mu\nu}_{{\bf Q}}\left({\bf p}\right)=\sum_{i}\left[{{\cal M}^{i}}\right]^{\mu\nu}\Phi_{{\bf Q}}^{i}\left({\bf p}\right).$ (12)
@@
The fluctuation amplitudes $\Phi_{\bf Q}$ can be related to the $a_{n,{\bf Q}}$ eigenmodes by the transformation [see Eqs. (24),(25)],
$a_{n,{\bf Q}}=\left[T^{-1}_{\bf Q}\Phi_{\bf Q}\right]_{n}.$ (21)
@@ III.2 Parametric term
$\delta{\cal H}=\frac{1}{2}\sum_{{\bf Q}}\begin{pmatrix}a_{{\bf Q}}^{\dagger}&a_{-{\bf Q}}\end{pmatrix}\delta{\Lambda}_{\bf Q}\begin{pmatrix}a_{{\bf Q}}\\ a_{-{\bf Q}}^{\dagger}\end{pmatrix}$, (22)
$\delta{\Lambda}_{\bf Q}=\begin{pmatrix}\delta{\cal E}_{{\bf Q}}&\delta{\cal Z}_{{\bf Q}}\\ \delta{\cal Z}_{{\bf Q}}^{\dagger}&\delta{\cal E}_{{\bf Q}}\end{pmatrix}$, (23)
@@
We stress that the transformation matrix $T_{\bf Q}$ is generally not unitary. Finally, $\delta{\cal H}$ can be read off as,